A REMARK ON THE STRICT CONVEXITY FOR SEMINORMS II

by Ivan Franić, Zagreb, Croatia

Abstract. We give an equivalent characterization of the strict convexity for seminorms in the sense of C.Diminie and A.White by using a form of a duality mapping. Mathematics subject classification: 46A99 Keywords: strict convexity, seminorms.